OpenStudy (anonymous):

Problem for you #4 : A number ends with 7. Find out the smallest number such that this number becomes 5 times if 7 is shifted to the 1st position. Good Luck!!

OpenStudy (anonymous):

How do you define shifting? ABC7 -> 7ABC ?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

and 7ABC =5 * ABC7, right?

OpenStudy (apoorvk):

Well that means the no. (original) ends with '57'.

OpenStudy (anonymous):

@FFM Yeah. Excpet that you dont know if its 4 digit. But otherwise yes.

OpenStudy (experimentx):

it's my wild guess that first digit must be 1

OpenStudy (apoorvk):

Hmm. So so so... The first digit of the no. has got to be '1'. So, the whole thing is like, suppose there are 'n' digits, then..

OpenStudy (apoorvk):

Yeah @experimentX your wild guess ain't so wild!

OpenStudy (experimentx):

lol ... second digit must be 4 or 5 ...

OpenStudy (apoorvk):

\[5 \times (10^n + ...... + 5 \times 10^2 + 7) = 7 \times 10^n + 10^{n-1} + ......+ 5\]

OpenStudy (anonymous):

* LHS last is 5 * 10

OpenStudy (apoorvk):

something like this^. Yeah and now I guess, it's "find the next digit" game. Yes SORRY. - LHS's second last term would be 5x10 and NOT 5x10^2 - thanks @siddhantsharan

OpenStudy (experimentx):

mos likely the second digit is 4 ... and the third digit it 2 or 3

OpenStudy (anonymous):

Apoorv. You got the problem with your way?

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=142857*5 ??

OpenStudy (anonymous):

@experimentX That is correct. Method?

OpenStudy (experimentx):

it's hard to explain 2*5 = 10 ... so first digit is cannot be greater than 1 5*1 = 5 we we need 2 carry ... which means 4 and 5

OpenStudy (apoorvk):

yeah, that explanation for '1' is alright to me.

OpenStudy (experimentx):

4 or 5 !! it's difficult to get 5 9*5 = 45 <-- so i choose 4

OpenStudy (anonymous):

How'd you know the number of digits involved. Or wait. You just proceeded till you got a 7?

OpenStudy (experimentx):

something like this and that ... more of it .. hit and trial and a bit of luck

OpenStudy (anonymous):

Haha. You always do get there. Nice going man!! :D:D

OpenStudy (experimentx):

lol ... not really ... these types of question you usually ask never favors me ... sometimes it get lucky that's all

OpenStudy (experimentx):

any way thanks for the question too

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (anonymous):

Well I have solved it.

OpenStudy (anonymous):

Not a bad problem, I have solved something like this 2 years back.

OpenStudy (anonymous):

HINT: \( \frac 1 7 \)

OpenStudy (anonymous):

142857

OpenStudy (anonymous):

Precisely. Knew you would. How did you do it?

OpenStudy (anonymous):

How'd you do ittttt??

OpenStudy (anonymous):

I knew this property. I had solved this problem before PE-52 http://projecteuler.net/problem=52